OpenStudy (anonymous):

Math Analysis: Simplify each expressioncos^3 y + cos y*sin^2 y

OpenStudy (campbell_st):

well cos(y)sin(y) = 1/2sin(2y) so you get cos^3(y) + 1/2sin(2y) thats one possibility an other is cos(y) (cos^2(y) + sin(y)) = cos(y)(1 - sin^2(y) + sin(y)) don't know if there is a lot else you can do

OpenStudy (anonymous):

the answer from the back of the book said cos y

OpenStudy (anonymous):

so how would I get that

OpenStudy (campbell_st):

no idea...

OpenStudy (anonymous):

...

OpenStudy (chris):

not sure that's correct, but I may be wrong

OpenStudy (chris):

(they are not the same)

OpenStudy (anonymous):

Shouldn't it be \[\cos^3 y + \cos y \, \,\sin^2 y\], because it would be easy to get cos y

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

whoops i forgot to put the ^2 on the sin heheh ^^"

OpenStudy (campbell_st):

lol.... now that makes sense its cos(y) (cos^2(y) + sin^2(y)) = cos(y) (1) since cos^2 + sin^2 = 1 answer cos(y)

OpenStudy (anonymous):

it says cos^3 y in the equation not cos^2 y o-o

OpenStudy (anonymous):

cos^3 y + cos Y*sin^2y

OpenStudy (campbell_st):

take cos(y) as a factor so cos(y) x cos^2(y) = cos^3(y)

OpenStudy (anonymous):

i still dont get how the answer is cos y o-o

OpenStudy (campbell_st):

find the common factor between cos^3(y) + cos(y)sin^2(y) it is cos(y) so you can factorise the expression to \[\cos(y) \times (\cos^2 (y) + \sin^2(y))\]

OpenStudy (anonymous):

cos y doesn't factor sin^2 y

OpenStudy (anonymous):

ohh wait nvm i see now lol i feel like a fool now xD Tyvm :3